OpenStudy (anonymous):

what is the two-decimal place approximation of square root of 31???? please help

OpenStudy (amistre64):

make a guess g refine the guess using:\[g_n=\frac{31+g^2}{2g}\] when gn and gn+1 have the same positional values for n.xx you got it

OpenStudy (amistre64):

lets say 6 is a good guess, since 6^2 is 36 (31+6^2)/(2*6) = 4.666 (31+4.666^2)/(2*4.666) = 5.654762 after a 2 or 3 runs we get to 5.567764363 as a constant result on the calculator

OpenStudy (anonymous):

thank you1

OpenStudy (amistre64):

lol, but since im using a calculator, we could have just done sqrt(31) = 5.567764363 ...